OpenStudy (anonymous):

What is the value of the sum 5 + 10 + 15 + ... + 95 + 100?

OpenStudy (anonymous):

1050

OpenStudy (anonymous):

Is there an equation or formula to solve?

OpenStudy (anonymous):

\[S_n = \frac{n}{2} (\text{First Term} + \text{Last Term})\] It's an Arithematic Progression, maybe you know that already.

OpenStudy (anonymous):

Arithmetic*

OpenStudy (anonymous):

Ah! Thank you. Yes I just could not think of it for nothing!

OpenStudy (anonymous):

No Problem

myininaya (myininaya):

\[\sum_{i=1}^{20}5i=5\sum_{i=1}^{20}i=5 \cdot \frac{20(20+1)}{2}=5 \cdot 10(20+1)=50(21)=1050\]

OpenStudy (karatechopper):

i dont understand through ishaans formula. and myins is more than confusing!

myininaya (myininaya):

lol what

myininaya (myininaya):

no not confusing we have 5(1+2+3+...19+20)

myininaya (myininaya):

\[=5 \cdot \frac{20(20+1)}{2}\]

OpenStudy (karatechopper):

what is n

myininaya (myininaya):

\[\sum_{i=1}^{k} i=\frac{k(k+1)}{2}\]

OpenStudy (karatechopper):

is n 5?

myininaya (myininaya):

n is the number of terms in ishaan's expression

OpenStudy (karatechopper):

uhh?

OpenStudy (karatechopper):

o i see

OpenStudy (karatechopper):

tht would be 5/2(5+100)

myininaya (myininaya):

i bet satellite is fixing to say something to put down my way

OpenStudy (anonymous):

\[5\sum_{k=1}^{20}k=5\times \frac{20\times 21}{2}\] \[5\times 10\times 21\]

myininaya (myininaya):

lol

OpenStudy (anonymous):

no no i like it. what else would you do?

myininaya (myininaya):

ishaan's way

OpenStudy (karatechopper):

uhhh

myininaya (myininaya):

our his expression works great too

myininaya (myininaya):

or*

OpenStudy (karatechopper):

satellite i have no clue what tht funky llooking E stands for

OpenStudy (anonymous):

sum

myininaya (myininaya):

chopper it means we are adding a whole bunch and we don't want to write down every term of the finite series

OpenStudy (anonymous):

but what is that picture? andy warhol painting portraits of tourists at the beach?

OpenStudy (karatechopper):

i can make tht^

myininaya (myininaya):

cazil after you seen the E looking thing its actually called sigma

myininaya (myininaya):

have you* i can't type

myininaya (myininaya):

example: \[\sum_{i=1}^{6}f(x_i)=f(x_1)+f(x_2)+f(x_3)+f(x_4)+f(x_5)+f(x_6)\]

OpenStudy (anonymous):

I have seen it before. But am not very familiar with it yet. The arithmetic progression I know though.

myininaya (myininaya):

it comes very useful when you have infinite series

myininaya (myininaya):

you will see it more if you do more mathematics

OpenStudy (anonymous):

Did the formula help? I dislike memorizing formulas!

OpenStudy (karatechopper):

myin did u delete my pic?

OpenStudy (karatechopper):

it was fine if u did

myininaya (myininaya):

Maybe. I was trying to stick to the topic. Sorry.

OpenStudy (karatechopper):

it was fine! i was just joking around

myininaya (myininaya):

I know.

OpenStudy (karatechopper):

:) i still dont understand the whole ishaan formula tho..

myininaya (myininaya):

which part?

myininaya (myininaya):

you get the n part right?

OpenStudy (karatechopper):

the formula it self no the n part if 5=n then it still needs to be divided by 2 right?

myininaya (myininaya):

well n=20 we have 20 terms ishaan's formula we have \[20 \cdot \frac{5+100}{2}=10(105)=1050\]

OpenStudy (karatechopper):

but it says n/2

myininaya (myininaya):

yes i divided by 2

OpenStudy (karatechopper):

no i mean that 20 should be divided by 2

OpenStudy (anonymous):

kc i... generally don't recommend cramming, but since it's a special case, i advise you to learn the formula and you can deduce and understand it later.

OpenStudy (karatechopper):

no i am trying to make a point here. im learning

OpenStudy (karatechopper):

@ myin, u wrote (5+100)/2 but shouldnt it be 20/2?

OpenStudy (anonymous):

no have a look at the formula again \[S_n = \frac{n}{2} (\text{First Term} + \text{Last Term})\] n is the number of terms of a sequence

myininaya (myininaya):

\[\frac{n}{2}(first term+last term)=n \frac{first term+last term}{2}\]

myininaya (myininaya):

same thing

OpenStudy (karatechopper):

ok

OpenStudy (karatechopper):

so the 20/2(5+100) 10(105) 1050?

myininaya (myininaya):

yes

OpenStudy (karatechopper):

YAY! i learned something new! thank you!

OpenStudy (anonymous):

n in this particular sequence of 5, 10, 15 ... 100 is 100 = 5 + (n-1)5 95 = (n-1)5 19 + 1 = n => n =20 apply the formula \[\S_{20} = \frac{20}{2}* (5 + 100) = \frac{5 + 100}{2}\times 20\]

OpenStudy (karatechopper):

i see tht bhaiya i understnad it now